OpenStudy (anonymous):

how to find d 0.25(d-12)=4

OpenStudy (anonymous):

4/.25=d-12

OpenStudy (anonymous):

Distribute first bro. So, any term on the outside of a bracket, I think of it like its gotta go into the bracket and shake hands with everyone in there... kind of like a party you know? So, 0.25 x d = 0.25d. So then it's gotta go in there and shake hands with -12. So, 0.25 x -12 = -3 So now you have: 0.25d - 3 = 4 You have to solve for d Bring that 3 over to the other side, 0.25d = 4+3 0.25d = 7 divide 7 by 0.25 to get d d = 7/0.25 d = 28.

OpenStudy (anonymous):

or you can divide the 4 by the .25 and not worry about distribution. 4/.25=16 16=d-12 28=d

OpenStudy (anonymous):

Yeah, I know, but it seems like he didn't know how to even begin the problem the simple way... so for future reference when there's more complicated problems involving brackets, it might be a good concept to know.

OpenStudy (anonymous):

thanks